# Characterizing graph-nonedge pairs with single interval Cayley configuration spaces in 3-dimensions

### v1

@@ B.5. Proof of Proposition 5[$(G,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected]
Note that $L\cup xw$ is $3$-connected and $L$ is one of its $3$-connected spanning proper subgraphs. If we can show that $L\cup xw$ is $3$-flattenable, then the lemma follows from Lemma 20. Consequently, assume that $L\cup xw$ is not $3$-flattenable, and hence has a forbidden minor. By Lemma 18, $xw$ is contracted in every such minor. Hence, Proposition 1 shows that $(L,xw)$ is a winged graph pair. Therefore, either Proposition 3, if some $f$-winged graph minor of $L\cup xw$ has singleton wing tips, or Lemma 12, otherwise, shows that $L$ has a separator of size two. However, this contradicts the fact that $L$ is $3$-connected. Thus, $L\cup xw$ has no forbidden minor, and so the proof is complete. ∎
@@
Let $\{x_{i}y_{i}\}$ be the ordered set of pins of the chained $f$-winged graph minor $[G\cup f]$. Consider the links $L_{0}$, containing $u$ and $x_{1}y_{1}$, and $L_{1}$, containing $x_{1}y_{1}$ and $x_{2}y_{2}$, of $[G\cup f]$, and let $L=L_{0}\cup L_{1}$. By Lemma 19, we can assume wlog that $L_{0}$ and $L_{1}$ are both $3$-connected. In this case, we will show that $(G,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected. Since $[G\cup f]$ has singleton wing tips and its pins are edges of $G$, it suffices to show that $(L,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected. By definition, we must show that the CCS $\Omega^{3}_{\{ux_{2},uy_{2}\}}(L,\ell)$ is connected for any squared edge-length map $\ell$. Consider the CS $\mathcal{C}^{3}(L,\ell)$ and its set of connected components $\{X_{i}\}$. Let $Y_{i}$ be the image of $X_{i}$ under the Cayley map $\phi_{ux_{2},uy_{2}}$. Observe that $\Omega^{3}_{\{ux_{2},uy_{2}\}}(L,\ell)$ is the union of the sets $Y_{i}$. Furthermore, each $Y_{i}$ is connected since $X_{i}$ is connected and $\phi_{ux_{2},uy_{2}}$ is continuous. Therefore, we complete the proof by showing that $Y_{i}\cap Y_{j}\neq\emptyset$ for any $i$ and $j$.
@@
Let $Y_{i}(ux_{2})$ and $Y_{j}(ux_{2})$ be the coordinate projections of $Y_{i}$ and $Y_{j}$ onto $ux_{2}$, respectively. Also, let $Y_{i}(uy_{2})$ and $Y_{j}(uy_{2})$ be the coordinate projections of $Y_{i}$ and $Y_{j}$ onto $uy_{2}$, respectively. Since $Y_{i}$ is connected, it contains a continuous curve $\tau_{i}$ whose coordinate projections are $Y_{i}(ux_{2})$ and $Y_{i}(uy_{2})$. Similarly, $Y_{j}$ contains a continuous curve $\tau_{j}$ whose coordinate projections are $Y_{j}(ux_{2})$ and $Y_{j}(uy_{2})$. Additionally, Lemma 22 can be used to show that $Y_{i}(ux_{2})=Y_{j}(ux_{2})$ and $Y_{i}(uy_{2})=Y_{j}(uy_{2})$. Therefore, $\tau_{i}$ and $\tau_{j}$ are contained in $Y_{i}(ux_{2})\times Y_{i}(uy_{2})$. Thus, the Intermediate Value theorem shows that these curves share some point, and so $Y_{i}\cap Y_{j}\neq\emptyset$. This completes the proof. ∎
Appendix C Proofs of the forward direction propositions in Section 4
C.1. Proof of Proposition 6 [No Type (2) edge implies non-$3$-SIP map]
The proof of Proposition 6 requires Lemmas 23 and 25, below.
Lemma 23 (No Type (2) edge properties).
@@ C.1. Proof of Proposition 6 [No Type (2) edge implies non-$3$-SIP map]
Let $(G,f)$ be a minimal graph-nonedge pair, where $f=uv$, $[G\cup f]$ be an $f$-preserving forbidden minor, and consider the subgraph $H=[G\cup f]\setminus\{[u],[v]\}$. If $G$ has no Type (2) edge, then the following statements are true:
  1. (i)
  2. (ii)
    $|[w]^{-1}|=1$ for any vertex $[w]\in V(H)$,
  3. (iii)
  4. (iv)
    the sets $[u]^{-1}$ and $[v]^{-1}$ each have cardinality at most $2$.
Proof.
For Statement (ii), consider any vertex $[w]\in V(H)$. If $|[w]^{-1}|>1$, then there exists an edge of $G$ that is contracted in $[G\cup f]$ and that does not share an endpoint with $f$. However, this contradicts Statement (i), and so we have $|[w]^{-1}|=1$.
For Statement (iii), if $([u]^{-1}\cup[v]^{-1})\setminus\{u,v\}$ is empty, then we are done. Otherwise, let $x$ be any vertex in this set and assume wlog that $x\in[u]^{-1}$. Since $x\notin\{u,v\}$, there exists an edge $xy$ of $G$ that is contracted in $[G\cup f]$. Statement (i) tells us that $u$ is the only neighbor of $x$ in $[x]^{-1}\setminus\{u,v\}$. Furthermore, $xu$ is clearly not of Type (1), and so it must be of Type (3). Hence, $xv$ is an edge of $G\cup f$. It remains to show that $x$ neighbors at least two vertices in $H^{-1}$.
Assume that $x$ does not neighbor any vertex in $H^{-1}$. Since $G\cup f$ is an atom, by assumption, $x$ neighbors some vertex $y$ other than $u$ or $v$. By assumption and the discussion above, we have $y\in[v]^{-1}\setminus\{v\}$. Let $J$ be the subgraph of $G\cup f$ induced by $[u]^{-1}$. By the discussion above, $J\setminus x$ is connected and contains $u$. Hence, the minor obtained via the vertex exchange (see Section B.1) from $[u]^{-1}$ to $[v]^{-1}$ that exchanges $x$ and fixes $J\setminus x$ is an $f$-preserving forbidden minor in which $xy$ is contracted. However, since neither $x$ nor $y$ is an endpoint of $f$, this contradicts Statement (i), and so $x$ must neighbor some vertex $w$ in $H^{-1}$. If $w$ is the only vertex in $H^{-1}$ that $x$ neighbors, then $xw$ is contracted in the minor obtained via the vertex exchange from $[u]^{-1}$ to $[w]^{-1}$ that exchanges $x$ and fixes $J\setminus x$, and this minor either is or has an $f$-preserving forbidden minor. Therefore, we arrive at a similar contradiction, and so $x$ neighbors at least two vertices in $H^{-1}$.
Lemma 25 is proved using Lemma 24, below.
@@
Let the vertex set of $G$ be $\{v_{i}\}$, where $H$ does not contain $v_{1}$, $f=v_{1}v_{2}$, and $f^{\prime}=v_{3}v_{4}$. Also, let $a=\ell_{1}(v_{2}v_{3})$, $b=\ell_{1}(v_{2}v_{4})$, $c=\ell_{1}(v_{3}v_{4})$, and $d=\ell_{2}(v_{3}v_{4})$. Wlog, assume that $c<d$. Set $\ell_{1}^{\prime}(v_{1}v_{3})$ and $\ell_{1}^{\prime}(v_{1}v_{4})$ to be positive numbers such that $c=|\ell_{1}^{\prime}(v_{1}v_{3})-\ell_{1}^{\prime}(v_{1}v_{4})|$, $d=\ell_{1}^{\prime}(v_{1}v_{3})+\ell_{1}^{\prime}(v_{1}v_{4})$, and $\ell_{1}^{\prime}(v_{1}v_{3})$ is not equal to $a$ or $b$. It is easy to see that Conditions (i)-(iii) are satisfied, and so the lemma is proved. ∎

### v2

@@ 4.2   Proofs of Tool Theorems 5 and 6 concerning partial 3-trees and covering maps
Since $G\cup F$ is $3$-flattenable, Theorem 3 shows that the CCS is convex. Hence, each point $y\in Y$ is not on any boundary of the CCS. This implies that each realization in $\phi_{F}^{-1}(y)$ is contained in a $\delta$-oriented CS for some $\delta\in\Delta$. Since $G\cup F$ is a $3$-tree, it follows that, for each $\delta\in\Delta$, the $\delta$-oriented CS contains some realization in $\phi_{F}^{-1}(y)$, which proves (i). Next, (ii) follows easily from the fact that the CCS is the semi-algebraic set defined by a set of tetrahedral inequalities: one for each $K_{4}$ subgraph of $G\cup F$ and subject to the edge-length map $\ell$. ∎
Lemma 17 (Oriented CSs are connected).
Let $G$ be a graph with $F$ being a subset of its nonedges such that $G\cup F$ is a $3$-tree, let $\ell$ be a squared edge-length map, and let $\phi_{F}$ be the Cayley map from the CS $\mathcal{C}^{3}(G,\ell)$ to the CCS $\Omega^{3}_{F}(G,\ell)$. Also, consider the set $A=\{A_{1},\dots,A_{m}\}$ containing each $K_{4}$ subgraph of $G\cup F$ that maps to a non-coplanar set of points under some realization in the CS. If the set $A$ is empty, then the CS is connected; otherwise, for each $\delta\in\Delta$, the $\delta$-oriented CS is connected.
The proof of Lemma 17 is given in A.3.
Proof of Theorem 6.
We must show that $(G,F)$ has the $3$-covering map property whenever $G\cup F$ is a partial $3$-tree. It suffices to consider the case where $G\cup F$ is a $3$-tree. Let $\ell$ be a squared edge-length map and let $\phi_{F}$ be the Cayley map from the CS $\mathcal{C}^{3}(G,\ell)$ to the CCS $\Omega^{3}_{F}(G,\ell)$. We must show that the image of each connected component of the CS under $\phi_{F}$ is the CCS. Let $A=\{A_{1},\dots,A_{m}\}$ be the set containing each $K_{4}$ subgraph $A_{i}$ of $G\cup F$ that maps to a non-coplanar set of points under some realization in the CS. If $A$ is empty, then Lemma 17(i) shows that the CS is connected, and so the theorem is immediate. Otherwise, Lemma 16(ii) and Lemma 17(ii) show that each connected component in the CS contains a $\delta$-oriented CS for some $\delta\in\Delta$. Lemma 16(i) then states that the image of this $\delta$-oriented CS under $\phi_{F}$ is $Y$. The theorem now follows since $\phi_{F}$ is a closed map. ∎
Now we are ready to state and prove Proposition 5.
Proposition 5 ($(G,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected).
@@
Consider the links $L_{0}$, containing $u$ and $x_{1}y_{1}$, and $L_{1}$, containing $x_{1}y_{1}$ and $x_{2}y_{2}$, of $[G\cup f]$, and let $L=L_{0}\cup L_{1}$. By Lemma 12, we can assume wlog that $L_{0}$ and $L_{1}$ are both $3$-connected. In this case, we will show that $(G,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected. Since $[G\cup f]$ has singleton wing tips and its pins are edges of $G$, it suffices to show that $(L,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected. By definition, we must show that the CCS $\Omega^{3}_{\{ux_{2},uy_{2}\}}(L,\ell)$ is connected for any squared edge-length map $\ell$. Consider the CS $\mathcal{C}^{3}(L,\ell)$ and its set of connected components $\{X_{i}\}$. Let $Y_{i}$ be the image of $X_{i}$ under the Cayley map $\phi_{ux_{2},uy_{2}}$. Observe that $\Omega^{3}_{\{ux_{2},uy_{2}\}}(L,\ell)$ is the union of the sets $Y_{i}$. Furthermore, each $Y_{i}$ is connected since $X_{i}$ is connected and $\phi_{ux_{2},uy_{2}}$ is continuous. Therefore, we complete the proof by showing that $Y_{i}\cap Y_{j}\neq\emptyset$ for any $i$ and $j$.
@@
Let $Y_{i}(ux_{2})$ and $Y_{j}(ux_{2})$ be the coordinate projections of $Y_{i}$ and $Y_{j}$, respectively, onto $ux_{2}$. Also, let $Y_{i}(uy_{2})$ and $Y_{j}(uy_{2})$ be the coordinate projections of $Y_{i}$ and $Y_{j}$, respectively, onto $uy_{2}$. Since $Y_{i}$ is connected, it contains a continuous curve $\tau_{i}$ whose coordinate projections are $Y_{i}(ux_{2})$ and $Y_{i}(uy_{2})$. Similarly, $Y_{j}$ contains a continuous curve $\tau_{j}$ whose coordinate projections are $Y_{j}(ux_{2})$ and $Y_{j}(uy_{2})$. Additionally, Lemma 13 can be used to show that $Y_{i}(ux_{2})=Y_{j}(ux_{2})$ and $Y_{i}(uy_{2})=Y_{j}(uy_{2})$. Therefore, $\tau_{i}$ and $\tau_{j}$ are contained in $Y_{i}(ux_{2})\times Y_{i}(uy_{2})$. Thus, the Intermediate Value Theorem shows that these curves share some point, and so $Y_{i}\cap Y_{j}\neq\emptyset$. This completes the proof. ∎
Finally, we can complete the proof of the converse direction of Theorem 4.
Proof of the converse direction of Theorem 4.
Let $(G,f)$ be a graph-nonedge pair such that no atom of $G\cup f$ that contains $f$ has an $f$-preserving forbidden minor. We will show that $(G,f)$ has the $3$-SIP by induction of the number $n$ of vertices in $G$. When $n\leq 5$, our assumption implies that $G\cup f$ has no forbidden minor, and so Theorem 3 shows that $(G,f)$ has the $3$-SIP. Next, assume that the converse direction of Theorem 4 is true when $n\leq k$, for any integer $k\geq 5$, and we will prove it when $n=k+1$. If $G\cup f$ has no forbidden minor, then Theorem 3 shows that $(G,f)$ has the $3$-SIP. Otherwise, Corollary 3 allows us to assume wlog that $G\cup f$ is an atom. Since atoms are $2$-connected, Proposition 1 shows that $(G,f)$ is a winged graph pair. Consequently, Propositions 2, 3, and 4 allow us to assume that $(G,f)$ has a chained $f$-winged graph minor $[G\cup f]$ with singleton wing tips. Hence, let $\{x_{i}y_{i}\}$ be the ordered set of pins of $[G\cup f]$. Proposition 5 shows that $(G,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected, where $f=uv$. Therefore, by Lemma 2, it suffices to show that $(G\cup\{ux_{2},uy_{2}\},f)$ has the $3$-SIP, which wedo next.
Observe that for some vertex $w\in\{x_{1},y_{1}\}$, $\{u,x_{2},y_{2}\}$ is a clique $wv$-separator of $G\cup\{f,ux_{2},uy_{2}\}$. Hence, by Lemma 3, it suffices to consider the $\{u,x_{2},y_{2}\}$-component $H$ of $G\cup\{f,ux_{2},uy_{2}\}$ that contains $f$ and show that $(H\setminus f,f)$ has the $3$-SIP. If this is not the case, then, since $|V(H)|<|V(G)|$, the inductive hypothesis states that $H$ has an $f$-preserving forbidden minor. However, since both $\{x_{1},y_{1}\}$ and $\{x_{2},y_{2}\}$ are minimal $uv$-separators of $G$, there exist paths in the $\{u,x_{2},y_{2}\}$-component of $G\cup f$ that contains $w$ between $w$ and each vertex in the set $\{u,x_{2},y_{2}\}$, and each path contains exactly one vertex in $\{u,x_{2},y_{2}\}$. This implies that some minor $[G\cup f]_{1}$ is $H$ with $[f]_{1}$ as $f$. Consequently, $G\cup f$ has an $f$-preserving forbidden minor, which is a contradiction. Thus, $(H\setminus f,f)$ has the $3$-SIP, and so the proof is complete. ∎
5   Proof of the forward direction of main Theorem 4
We first outline the argument before filling in the details. See Figure 18 for a proof diagram. We use the same notational convention as in the converse direction of the proof of Theorem 4: since the notation $(G,f)$ is used to refer to general graph-nonedge pairs in the inductive and other steps of the proof, we designate the notation $(G_{\star},f_{\star})$ throughout the proof for the particular pair in Theorem 4. Assume some atom of $G_{\star}\cup f_{\star}$ contains $f_{\star}$ and has an $f_{\star}$-preserving forbidden minor $[G_{\star}\cup f_{\star}]$, which we can ensure is an induced minor. We show that $(G_{\star},f_{\star})$ does not have the $3$-SIP. Proposition 6 handles the two cases where $G_{\star}\cup f_{\star}$ is one of the forbidden minors $K_{5}$ or $K_{2,2,2}$. Otherwise, if $f_{\star}$ is retained in the induced forbidden minor $[(G_{\star},f_{\star})]$, then combining this proposition with Lemma 1 completes the proof. Thus most of the proof focuses on the case where $f_{\star}$ is preserved but not retained, i.e., doubled, in a forbidden minor. In this case, Lemma 1 does not apply, and so the main obstacle is propagating the non-$3$-SIP certificate from the forbidden minor edges to $f_{\star}$.
Figure 16: All minimal graph-nonedge pairs with no Type (2) edge. See the discussion, Example 6, Lemma 19, and Proposition 6 in this section.
Take for example $(G_{\star},f_{\star})$ to be the pair in either Figure 16(c) or 16(d), where exactly one edge $e$ is contracted to reach $[G_{\star}\cup f_{\star}]$ and $[f_{\star}]^{-1}$ contains exactly one edge $f^{\prime}$ other than $f_{\star}$. Proposition 6 provides a squared edge-length map $\ell_{1}$ for $G_{\star}\setminus f^{\prime}$ with $\ell_{1}(e)=0$ and for which the CCS $\Omega^{3}_{\{f_{\star},f^{\prime}\}}(G_{\star}\setminus f^{\prime},\ell_{1})$ is two disjoint intervals. See the plane where $\ell(e)=0$ in the schematic illustration in Figure 17. Observe that any section of the CCS that fixes the length of $f^{\prime}$ contains exactly one point, and so $\ell_{1}$ cannot witness the non-$3$-SIP property of $f_{\star}$. This property on sections holds precisely because $f_{\star}$ is doubled in $[G_{\star}\cup f_{\star}]$. We argue that there exists a squared edge-length map $\ell_{2}$ for $G_{\star}\setminus f^{\prime}$ that agrees with $\ell_{1}$, except on $e$ which it maps to some real number $\epsilon>0$, such that some section of the CCS $\Omega^{3}_{\{f_{\star},f^{\prime}\}}(G_{\star}\setminus f^{\prime},\ell_{2})$ that fixes the length of $f^{\prime}$ is the union of disjoint intervals. This section is schematically illustrated as the intersection of the red fiber and the solid disks in Figure 17.
Thus, letting $\ell^{\prime}$ be the squared edge-length map for $G_{\star}$ that agrees with $\ell_{2}$ and sets $\ell^{\prime}(f^{\prime})$ to be the length of $f^{\prime}$ in this section demonstrates that the CCS $\Omega^{3}_{f_{\star}}(G_{\star},\ell^{\prime})$ is two disjoint intervals, which would complete the proof of the forward direction of the main Theorem 4.
$J$$[H]$$\ell(f^{\prime})$$\ell(e)$$\ell(f_{\star})$$\ell(e)=0$$\ell(e)=\epsilon$
Figure 17: See the proof sketch at the beginning of this section. Shown are two solid cones (drawn hollow for clarity) that schematically illustrate the CCS $\Omega^{3}_{\{f_{\star},f^{\prime},e\}}(G_{\star}\setminus\{f^{\prime},e\},\ell)$. Two sections of this CCS that fix the length $\ell(e)$ for $e$ to be 00 and $\epsilon>0$, respectively, are shown in dark-blue: the former consists of two line-segments and the latter consists of two ellipses. These sections schematically illustrate the CCSs $\Omega^{3}_{\{f_{\star},f^{\prime}\}}(G_{\star}\setminus f^{\prime},\ell_{1})$ and $\Omega^{3}_{\{f_{\star},f^{\prime}\}}(G_{\star}\setminus f^{\prime},\ell_{2})$, respectively. All sections of $\Omega^{3}_{\{f_{\star},f^{\prime}\}}(G_{\star}\setminus f^{\prime},\ell_{1})$ that fix a length for $f^{\prime}$ consist of a single point, as illustrated by the intersection of the red fiber with the CCS. Some section of $\Omega^{3}_{\{f_{\star},f^{\prime}\}}(G_{\star}\setminus f^{\prime},\ell_{2})$ that fixes a length for $f^{\prime}$ consists of two disjoint intervals, as illustrated by the intersection of the red fiber with the CCS. This section is the CCS $\Omega^{3}_{f_{\star}}(G_{\star},\ell^{\prime})$, which demonstrates that $(G_{\star},f_{\star})$ does not have the $3$-SIP.
While this proof sketch is inherently geometric, we lack the tools to prove it directly using properties of the Euclidean distance cone. Instead, we devise a lengthy combinatorial argument that reduces the problem to solving a few elementary geometry problems. First, Lemmas 1 and 3 (repeatedly used tools of independent interest stated in Section 2) allow us to assume $(G_{\star},f_{\star})$ is minimal as defined below. Then Theorem 7 establishes that minimal pairs do not have $3$-SIP, which then completes the forward direction of 4.
Definition 2 (Types of edges of a graph-nonedge pair).
Given a graph-nonedge pair $(G,f)$, the edges of $G$ can be partitioned into the following four classes. Let $[G\cup f]$ be the minor obtained by contracting any one edge of $G$. Then, the contracted edge is of Type:
  1. 1.
    if $[G\cup f]$ has no $[f]$-preserving forbidden minor,
  2. 2.
    if $[G\cup f]$ has some $[f]$-preserving forbidden minor, but no atom of $[G\cup f]$ that contains $[f]$ has such a minor,
  3. 3.
  4. 4.
    if none of the above apply, in which case the edge is called reducing.
Definition 3 (Minimal pairs).
@@ 5   Proof of the forward direction of main Theorem 4
A graph-nonedge pair $(G,f)$ is minimal if $G\cup f$ is an atom that has an $f$-preserving forbidden minor and $G$ has no reducing edge.
Example 6.
A graph-nonedge pair $(G,f)$ is not minimal if it is the pair (i) in Figure 4(a) even though $G\cup f$ has an $f$-preserving forbidden minor, since $G\cup f$ is not an atom, or (ii) on the bottom-right of Figure 4(c) even though $G\cup f$ is an atom that has an $f$-preserving forbidden minor, since the edge whose contraction yields an $f$-retaining $K_{5}$ minor is reducing. $(G,f)$ is minimal if it is any of the pairs in Figure 16 or if $G\cup f$ is any of the graphs in Figures 7(a), 7(b), 7(d), 7(f), 7(g), or 7(i). ∎
Remark 3.
Notice that our definition of minimal pair does not prevent one minimal pair $(G,f)$ from properly containing another minimal pair $(G^{\prime},f^{\prime})$ where $G^{\prime}$ is a proper induced subgraph of $G$. For example, this is true if $G\cup f$ is the graph in Figure 7(b) and $G^{\prime}\cup f^{\prime}$ is the graph in Figure 7(a); or $G\cup f$ is the graph in Figure 7(d) and $G^{\prime}\cup f^{\prime}$ is the graph in Figure 7(b) with $f$ and $f^{\prime}$ being solid green.
Next, Theorem 7, below, demonstrates that no minimal pair has the $3$-SIP. A squared edge-length map $\ell$ is a proper non-$d$-SIP map for $(G,f)$ if $\ell$ maps each edge of $G$ to a positive real number and the CCS $\Omega^{3}_{f}(G,\ell)$ contains exactly two distinct positive real numbers. For example, a squared edge-length map for the pair in Figure 1(d) assigns length $1$ to each in one $K_{4}$ subgraph of $G$ and length $2$ to the remaining three edges of $G$.
Forward direction of Theorem 4Theorem 7Proposition 8Lemma 34Lemma 33Lemmas 54,55,61,66,67Lemma 53Lemmas 52,61,63,65,68Lemma 30Lemmas 63,66,68Lemma 29Lemmas 4,5,62Lemma 28Lemma 27Lemma 26Lemma 42Lemma 43Lemma 44Lemma 25Lemma 32Lemma 31Lemmas 49-51Lemma 48Lemma 47Lemma 46 Lemma 45Lemma 68Proposition 7Lemma 23Lemma 22Lemma 60Lemma 21Proposition 6Lemma 20Lemma 19Lemma 18Proposition 9Lemma 35Lemma 56Proposition 10Lemma 24Lemma 36Lemmas 4,5,57-59,66,67
Figure 18: Proof outline for the forward direction of Theorem 4. See the discussion in this section.
Theorem 7 (Every minimal pair has a proper non-$3$-SIP map).
Next, we outline the proof of Theorem 7 using Propositions 6-10 below, which is by induction on the number $n$ of vertices in $G_{\star}$. The base case, where $G_{\star}\cup f_{\star}$ is $K_{5}$, is handled by Proposition 6. In the inductive step, this proposition lists all minimal graph-nonedge pairs with no Type (2) edge, shown in Figure 16, and shows that they have proper non-$3$-SIP maps. A minimal pair whose graph has a Type (2) has the property that $f_{\star}$ is doubled in any forbidden minor in which it is preserved. This is the same predicament as in the above geometric proof sketch, which we now handle by exploiting the structure of minimal pairs with Type (2) edges. In particular, we show that such a pair contains nested smaller minimal pairs (containing different nonedges), the smallest being one of the pairs in Figure 16. Hence, a natural first approach is to consider the set of all minimal pairs that properly contain exactly one minimal pair, and then proceed bottom-up. A subset of this set is shown in Figure 19; however, it is not clear how to produce the entire set. Instead, we proceed top-down as detailed in following example.
Figure 19: Minimal pairs (nonedge in red) that properly contain exactly one minimal pair. From left to right, the subgraphs of red and blue edges are the graphs in Figures 20(a), 20(c), 20(d), and 20(e), respectively. See the proof sketch at the beginning of this section.
Let $G_{\star}\cup f_{\star}$ be the graph in Figure 7(b) with $f_{\star}$ being the green edge, let $[G_{\star}\cup f_{\star}]$ be the minor shown in Figure 7(c), which is obtained by contracting a Type (2) edge; and let $C$ and $M$ be the purple and $K_{5}$ subgraphs of $[G_{\star}\cup f_{\star}]$, respectively. Note that $C$ is a CMS that separates $f_{\star}$ from some the forbidden minor $M$. Proposition 7 guarantees the existence of such a CMS. Let $H$ be the $C$-component of $[G_{\star}\cup f_{\star}]$ that contains $f^{\star}$, shown in the dashed black box. Consider the subgraphs $C^{-1}$ and $H^{-1}$ of $G_{\star}\cup f_{\star}$, and let $G^{\prime}=(G_{\star}\cup f_{\star})\setminus(H^{-1}\setminus C^{-1})$. Note that $G^{\prime}$ contains the smaller minimal pair in Figure 7(a). Choosing the CMS $C$ carefully, as we have, ensures that $H^{-1}$ is not a subgraph of any minimal pair properly contained in $(G_{\star},f_{\star})$. Figures 7(d) and 7(e) illustrate CMSs with and without this property. This very strong property forces a special structure on $H^{-1}$ described in Proposition 9: it is one of the graphs in Figures 20(a)-20(e). The proof of this proposition is highly technical and is given in Section 9. We also get Proposition 8, which states that $(G^{\prime},f^{\prime})$ has a proper non-$3$-SIP map $\ell^{\prime}$, where $f^{\prime}$ is the nonedge of $C^{-1}$ in our example. This argument has reduced our initial problem of showing $(G_{\star},f_{\star})$ does not have the $3$-SIP to solving elementary geometry problems about the graphs in Figures 20(a)-20(e). These problems are detailed and solved in Proposition 10, which extends $\ell^{\prime}$ to a proper non-$3$-SIP map $\ell$ for $(G_{\star},f_{\star})$. This completes the proof outline of Theorem 7.
Remark 4.
Figure 19 shows minimal pairs that properly contain exactly one minimal pair. Observe that all but one of the graphs in Figures 20(a)-20(e) appear as $H^{-1}$ in these pairs. We were not able to find an example of a minimal pair with $H^{-1}$ being the graph in Figure 20(b), which possibly means that this graph cannot be $H^{-1}$ for any minimal pair or that it only appears as $H^{-1}$ for sufficiently large minimal pairs. Our top-down proof covers both cases.
We now prove the above-mentioned Propositions 6-10, starting with Proposition 6, which lists all minimal pairs with no Type (2) edge and shows they each have a proper non-$3$-SIP map. The proof requires Lemmas 19 and 20, below. Lemma 19 is proved using Lemma 18, also below.
@@
Let the vertex set of $G$ be $\{v_{i}\}$, where $H$ does not contain $v_{1}$, $f=v_{1}v_{2}$, and $f^{\prime}=v_{3}v_{4}$. Set $\ell_{1}^{\prime}(v_{1}v_{3})$ and $\ell_{1}^{\prime}(v_{1}v_{4})$ to be positive numbers such that $\ell_{1}(v_{3}v_{4})=|\ell_{1}^{\prime}(v_{1}v_{3})-\ell_{1}^{\prime}(v_{1}v_{4})|$ and $\ell_{2}(v_{3}v_{4})=\ell_{1}^{\prime}(v_{1}v_{3})+\ell_{1}^{\prime}(v_{1}v_{4})$. By swapping the values of $\ell_{1}^{\prime}(v_{1}v_{3})$ and $\ell_{1}^{\prime}(v_{1}v_{4})$, we can additionally ensure that $\ell_{1}^{\prime}(v_{1}v_{3})\neq\ell_{1}(v_{2}v_{3})$. It is easy to see that Conditions (i)-(iii) are satisfied, and so the lemma is proved. ∎